### v1

# Challenges in applying the magnetic field circulation in Ampère-Maxwell’s law

Abstract
We investigate how students in introductory electromagnetism courses apply the idea behind the magnetic field circulation in the context of the Ampère-Maxwell law. For this purpose, we designed a metacognitive pencil-and-paper question, presented it to 65 students, and analysed their responses using phenomenography. We complemented our research by conducting interviews with 12 students. The results show that students tend to think that they can only use the magnetic field circulation to calculate magnetic fields if the curve used is symmetric, and that they do not take into account the shape of the magnetic field lines when applying it. We also find that some students believe that a variable electric field must cross the entire Amperian curve in order to apply Ampère-Maxwell’s law and find the magnetic field.
1 Introduction
Maxwell’s equations are the cornerstone of classical electromagnetic theory. In the realm of Physics Education Research (PER), there have been many studies of the conceptual difficulties that students face in applying them, with particular attention paid to Gauss’s law of the electric field [1, 2, 3, 4, 5, 6, 7, 8] and Faraday’s law [9, 10, 11, 12, 13], but less research on Ampère’s law and, in particular, Ampère-Maxwell’s law.
With regard to Ampère’s law, a significant percentage of students use reasoning based on memorisation and mechanical application of the information received in class and incorrectly consider that the magnetic field circulation can always be expressed as the product of the magnetic field and the length of the curve. They also think that the magnetic field used in the calculation of the line integral originates only from the currents crossing a surface bounded by the curve [1, 2]. Additionally, they assume that when the net current null, the magnetic field also vanishes [1, 2, 14]. Students often do not use the available information about the magnetic field to calculate the circulation, nor do they realise that the magnetic field circulation is a sum of elements $\vec{B}\cdot\overrightarrow{dl}$ along a closed curve [14]. They also confuse the concepts of circulation with magnetic field and circulation with flux [7, 8].
In relation to Ampère-Maxwell’s law, a recent work addresses whether students understand the limitations of Ampère’s law and how they understand and apply Ampère-Maxwell’s law [15] . The results of this study show that students struggle to recognise the limitations of Ampère’s law and to identify the current that crosses a surface bounded by a curve. It is also revealed that students have difficulty in recognising when a displacement current appears, as well as properly relating the magnetic field circulation to the rate at which the electric field varies.
This work is a continuation of research into students’ conceptual difficulties with Ampère-Maxwell’s law and is part of a wider project with the ultimate aim of developing a teaching-learning sequence. In this context, and considering the conceptual difficulties identified in the previous study, we have extended our original research to delve deeper into how students apply the magnetic field circulation in situations where they have to calculate it explicitly from the magnetic field. Our research question is:
  • How do students apply the magnetic field circulation in the context of Ampère-Maxwell’s law?
To address this research question, we designed a metacognitive pencil-and-paper question that led students to think about their own learning. This type of question has been widely recommended to investigate learning by understanding and active learning [16, 17]. The question presents a problem and a possible solution, so that the students must evaluate whether the solution is correct or not. To deepen the explanations received, we conducted semi-structured interviews with several groups of students. In the following sections, we describe the methodology used. We then present the results obtained, followed by a discussion and, finally, conclusions and implications for the teaching of Ampère-Maxwell’s law.
2 Experimental design
The question posed to the students shows a uniform electric field with magnitude increasing with time, confined to a cylindrical region of radius $R$, and states that the magnetic field is to be determined at a point $P$ located at a certain distance from the symmetry axis of the cylindrical region. The step-by-step solution proposed by a student to calculate the magnetic field as a function of the parameters of the problem is shown below. The proposed solution is correct, but it has the peculiarity that the closed curve chosen by the student exhibits a semicircle shape.
Presentation of the metacognitive question
Figure 1 shows a uniform electric field which increases with time and is confined to a cylindrical region of radius $R$. Agustina calculates the magnetic field at a point $P$ at a distance a from the centre of the cylindrical region.
Figure 1: Scheme presented to the students.
Analyse the steps that Agustina has taken to solve the problem and comment on each one, explaining why you agree or, if not, what changes you would make.
Solution: To find the magnetic field at $P$, I use a closed curve $C$ in the shape of a semicircle and apply Ampère-Maxwell’s law.
$\oint\vec{B}\cdot\overrightarrow{dl}=\mu_{0}I_{C}+\mu_{0}\varepsilon_{0}\frac{d\Phi_{E}}{dt}$ (1)
Since there are no conduction currents:
$\oint\vec{B}\cdot\overrightarrow{dl}=\mu_{0}\varepsilon_{0}\frac{d\Phi_{E}}{dt}$ (2)
where $d\Phi_{E}/dt$ is the change in electric field flux in the area enclosed by curve $C$.
Since the magnetic field circulation ( $\oint\vec{B}\cdot\overrightarrow{dl}$ ) along curve $C$ is $B\pi a$.
$B\pi a=\mu_{0}\varepsilon_{0}\frac{d\Phi_{E}}{dt}$ (3)
$B=\frac{\mu_{0}\varepsilon_{0}}{\pi a}\frac{d\Phi_{E}}{dt}$ (4)
The students were requested to analyse the previous solution and write a comment explaining whether they agree with it or, if not, what changes they would make. The solution proposed to the students is correct and is intended to challenge them in their self-reflection on the validity of the curve for calculating the magnetic field from the magnetic field circulation. Analysing the problem’s solution requires students to understand that:
  1. 1.
    The Ampère-Maxwell law is valid for any closed curve $C$, regardless of its symmetry.
  2. 2.
    To calculate the magnetic field circulation, it is necessary to know the shape of the field lines in the context of the defined curve $C$, since its definition is a sum of elements $\vec{B}\cdot\overrightarrow{dl}$.The Ampère-Maxwell law correlates the magnetic field circulation along the curve $C$ with the net current crossing a surface bounding it, without implying cause-effect relationships [18, 19].
3 Methodology
The question was answered by 65 undergraduate physics students in introductory electromagnetism courses at the University of the Basque Country (UPV/EHU) and the University of the Republic of Uruguay (UDELAR). These courses were taught by experienced university professors using standard teaching methods. The course duration was similar in both universities, lasting 15 weeks, with 4 hours of theory classes and an additional 2 hours per week dedicated to problem solving. The course content was based on the most recommended textbooks for introductory electromagnetism courses in the United States, Latin America and Spain [20].
The question was proposed to the students after they had received instruction on Ampère-Maxwell’s law and was analysed by two of the researchers using phenomenographic methodology [21]. This method made it possible to identify different categories of explanations that reflected the differences in students’ understanding. At the beginning of the analysis, each researcher evaluated 20 written tests. A consensus was then reached on the categories of understanding and explanation that emerged from the students’ responses. Each researcher then analysed the remaining tests individually. At the end of this process, the results obtained were compared, and the reliability of the analysis was assessed using Cohen’s kappa coefficient, which yielded a value of 0.95, indicating significant agreement [22].
To complement the study, we conducted semi-structured group interviews using the same question with three groups of two students and two groups of three students. Each interviews lasted approximately 20 minutes and in all the cases were conducted by the first author of this paper. The interviewer asked the students to discuss the question aloud and encouraged them to provide detailed explanations of their understanding using non-directive questions such as What do you mean by that?, Could you explain more?, Is there anything else you’d like to add about this step? All interviews were audio-recorded and transcribed. The extracts presented in the results section illustrate the sequence of interpretations and reasoning that characterise the processes of students’ development of situated cognition through problem solving. Fictitious names have been used in the presentation of results to ensure anonymity.
4 Results
From the phenomenographic analysis of students’ written responses, several explanatory categories emerge about how students reason when they engage in metacognitive reflection about solving the problem (see table 1).
Table 1: Answer categories and their percentages for question.
Category Description of the category %
A They agree with the analysis. 15.4
B They do not agree with the calculation of the magnetic field circulation. 15.4
C They do not agree with the choice of the curve. 20.0
D Other/Incoherent. 15.4
E No answer. 33.8
In category A we include students who justify their agreement with the solution presented. Let’s look at a representative example of this category:
“Yes, because in $C_{1}$, $\vec{B}\cdot\overrightarrow{dl}=0$ since they are orthogonal throughout $C_{1}$, and in $C_{2}$, $\vec{B}\cdot$ $\overrightarrow{dl}=Bdl$ since they are collinear.” (UDELAR-7, see Fig. 1)
Figure 2: Hand-drawn sketch given by student UDELAR-7.
Student UDELAR-7 divides the closed curve into two sections, $C_{1}$ and $C_{2}$, and correctly recognises that the contribution to the magnetic field circulation along $C_{1}$ is null. This indicates that, although they do not directly represent the magnetic field along $C_{1}$, they understand that the magnetic field is perpendicular to each element of the curve. On the other hand, the student observes that the magnetic field is collinear with section $C_{2}$. In this way, they demonstrate an adequate understanding of the shape of the magnetic field lines and correctly interpret and apply the concept of magnetic field circulation.
In category B we include answers that disagree with the calculation of the magnetic field circulation because they fail to recognise that the contribution of the magnetic field along the straight section of the curve is null. Let’s look at examples of category B:
“dl is $(2+\pi)a$ and the result is $B=\frac{\mu_{0}\varepsilon_{0}}{(2+\pi)a}\frac{d\Phi_{\mathrm{E}}}{dt}$.”(UPV-28)
“The induced field B is perpendicular to the field E at every point and its lines are closed. It is not perpendicular to the straight section of curve C, so it should have considered that section in the integral: $B(\pi a+2a)$.” (UDELAR-4)
Student UPV-28 believes that the magnetic field circulation should be calculated as the product of the magnetic field and the length of curve $C$. However, they do not recognise that the magnetic field along the straight section of curve $C$ does not contribute to the magnetic field circulation. On the other hand, student UDELAR-4 considers the direction of the magnetic field, acknowledging its perpendicularity to the electric field, but struggles to use this information effectively to calculate the magnetic field circulation, probably due to an incomplete understanding of circulation.
In the interviews we conducted, we found reasoning that gives us more clues about the different difficulties students exhibit in calculating the magnetic field circulation. In the following dialogue, two students exchange ideas about how to calculate the magnetic field circulation
Gustavo: To begin with, it does not take this part into account (in relation to the straight section), because $\pi a$ is only the curved part, not the diagonal.
Maite: Sure.
Gustavo: And if you took this diagonal into account (again with respect to the straight section), what would the field look like? Because the field and the dl are vectors and the angle has to be taken into account. So, I’m thinking what is this angle on this line, because if it were 90 you can cancel it, I think, if I remember correctly. I would say that they are circles (referring to magnetic field lines), they cannot be any other shape. It is impossible for it to be any other shape, in here they are circles (referring to the region where the electric field is), yes or yes. If they are circles, at any point here, you have a field vector like this and a dl like this, with an angle less than 90 (draw vectors $\vec{B}$ and $\overrightarrow{dl}$ along the straight section, see Fig. 3), so there is a magnetic field there and it doesn’t take it into account. That’s what’s wrong. She’s not taking that into account.
Interviewer: So what do you think is wrong with the solution of the problem?
Gustavo: That the integral has a value other than $B\pi a$.
Maite: Perhaps she is complicating her life by doing this. In other words, maybe if she had chosen the whole circle, it would have been much easier, because she does the circle and that’s it and everything is uniform there.
Gustavo: I think the same.
Figure 3: Hand-drawn sketch given by Gustavo during the interview.
Gustavo realises that the magnetic field circulation is a sum of the elementary terms $\vec{B}\cdot\overrightarrow{dl}$ and tries to calculate it by drawing magnetic field lines and $\overrightarrow{dl}$ segments. However, he does not realise that the contribution of the magnetic field to the circulation along the straight section is null because he has difficulty in correctly representing the magnetic field at different points along this segment.
In category C we group the answers that express disagreement with the choice of curve $C$ for the calculation of the magnetic field and support their arguments based on symmetry considerations. Let’s look at a few examples of category C:
“No, because in my case the integration curve would be $2\pi a$ and not $\pi a$.” (UPV-30)
“I think it is wrong, because the whole surface is not affected by the electric field, you have to consider the whole circle.” (UPV-5)
From student UPV-30’s answer, we can deduce that their disagreement with the solution of the question is due to the symmetry of the curve used, considering that a complete circle should be used. On the other hand, the UPV-5 student argues that a complete circle should be used so that the entire variable electric field crosses the surface bounded by the curve.
The students’ answers do not allow us to clearly identify the reasoning underlying their arguments. In the interviews, we found explanations that provide us with additional insights into the types of reasoning employed by students in this category. For example, during the analysis of the questions, one of the interviewees stated:
Lucía: The problem is the curve she has chosen. That is, Ampère-Maxwell’s law can be used to calculate magnetic fields, but the closed curve must have symmetry. The magnetic field must be the same at all points on the curve, and you also need to know the perimeter.
As we can see, Lucía clearly explains her reasoning. She states that the calculation of the magnetic field is only possible if the closed curve used is symmetrical.
In another interview, we identified an exchange between two students that is in line with what was expressed by the UPV-5 student. Below we present an extract from the interview in which the students exchange ideas about the shape of the closed curve.
Santiago: Surely (the question) is designed to induce an error, right? But I would consider the other half.
Interviewer: What do you mean by the other half?
Santiago: The part of the electric field that we forget.
Camila: Of course, you say that this is outside the curve (pointing to the area where there is an electric field that does not cross a surface bounded by $C$)
Santiago: Sure, it’s only enclosing half of it.
Camila: It should be a circle. Because it is the shape of the magnetic field generated by this variable electric field.
Santiago: Right. Because it covers the whole figure that has the electric field (referring to if a complete circle were used as a curve).
Camila: Sure, because there is a part of the electric field that you do not take into account, and if you do take it into account, well, we would have to see if it affects it or not.
Santiago: It affects, it affects, it affects. Maybe the parts that are farther away have less effect (referring to those that are farthest from point $P$), but they still have an effect on the magnetic field.
Santiago and Camila argue that the magnetic field at point $P$ is due to the entire electric field that crosses the surface bounded by the curve $C$. This idea leads them to conclude that the closed curve $C$ should be a complete circle so that the electric field completely crosses the surface bounded by the curve.
5 Discussion of results
When analysing the students’ answers to the research question, we found that around 15% adequately justify why they agree with all the steps in solving the question, showing an understanding of the application of magnetic field circulation in the context of Ampère-Maxwell’s law. We also observed that a similar percentage disagrees with the calculation of the magnetic field circulation, because they do not recognise that the magnetic field does not contribute to the circulation in the straight section. In this sense, we identified reasoning that could be related to the rote application of Ampère-Maxwell’s law, where students believe that the magnetic field circulation can always be expressed as the product between the magnetic field and the length of the closed curve. This type of reasoning has been documented in the literature in relation to Ampère’s law in different contexts [1, 2, 14].
Furthermore, we found that in some cases students do not consider the shape of the magnetic field lines when applying Ampère-Maxwell’s law, while in others they do not use the magnetic field information adequately to calculate the magnetic field circulation, nor do they recognise that the magnetic field circulation is a sum (see for example the first interview in section 4). These findings converge with those reported by Wallace and Chasteen [14] in the context of Ampère’s law.
We found that one in five students disagree with the choice of curve for calculating the circulation. Many of them believe that Ampère-Maxwell’s law can only be used to calculate magnetic fields if the curve used is symmetrical, which could be due to a functional fixation [23] associated with the most used procedure in class, which argues for symmetry to calculate magnetic fields using Ampère-Maxwell’s law. In addition, we found students who disagree with the shape of the curve used in the problem, considering that the variable electric field must cross the entire Amperian curve in order to be able to apply Ampère-Maxwell’s law and calculate the magnetic field. Our results suggest that students are unaware that Ampère-Maxwell’s law correlates the magnetic field circulation along a closed curve with the net current that crosses a surface bounding it.
This conceptual difficulty could be due to the belief that Ampère-Maxwell’s law establishes cause-effect relationships between the terms involved, in particular with the idea that currents (conduction and displacement) crossing a surface bounded by a curve must be the cause of the magnetic field computed in the line integral. This could lead them to argue that the curve must enclose the entire electric field. Another possible explanation is related to the belief that the magnetic field used in Ampère’s law is caused by enclosed currents [1, 2] and its origin could be in the way Maxwell’s equations are presented in most introductory physics textbooks, where they are described as cause-effect relationships between their different terms [20].
6 Conclusion and implication for teaching
In this article, we investigated how students in introductory electromagnetics courses apply magnetic field circulation in the context of the Ampère-Maxwell law. We designed a metacognitive pencil-and-paper question, presented it in writing to 65 students, and then analysed their answers using phenomenography. We complemented our research by conducting interviews with 12 students.
The circulation of a vector field is an essential part of the language in which Maxwell’s laws are expressed. Therefore, understanding and applying it correctly is essential to have a solid learning of the fundamental laws of electromagnetism. Our findings indicate that students still face significant difficulties when applying magnetic field circulation, some of which have been previously reported in relation to Ampère’s law. This suggests the complexity inherent in understanding and applying magnetic field circulation among students, regardless of the specific context.
We suggest that in order to promote a deeper understanding of magnetic field circulation in ths context, it is crucial to develop teaching-learning sequences that challenge students to calculate the magnetic field from this law, using unconventional curves and solving inverse problems where the magnetic fields cannot be determined from the line integral [7, 24]. This would allow students to reflect on the relationship between the symmetry of the problem, the fields and the circulation. In addition, it is essential to analyse magnetic field lines in different configurations involving conduction and displacement currents, and to apply the Ampère-Maxwell law to different surfaces bounded by the same curve [25, 26]. It is also worth mentioning the importance of recognising that the Ampère-Maxwell law does not provide information about the sources of the fields, nor does it imply causal relationships, but rather correlations between its different terms [18, 19].
Although the limitations of this study, focusing on a small sample of students, force us to be cautious about its generalisability, we believe that the consistency of the results with other studies on Maxwell’s laws allows us to draw some implications for the teaching of Ampère-Maxwell’s law and magnetic field circulation. In our next studies, we will investigate students’ conceptual difficulties that could arise from cause-and-effect interpretations of Maxwell’s equations, and we will use the results of this work, as well as those of previous research [20, 15], to design a teaching-learning sequence focused on Ampère-Maxwell’s law. This sequence will be implemented in introductory electromagnetism courses and evaluated to determine whether students overcome the conceptual difficulties.
Acknowledgment
[…]